OpenStudy (anonymous):

find the slope and y intercept of equation 2x-5y=3

OpenStudy (anonymous):

the slope is 2/5 y intercept is -3/5

OpenStudy (anonymous):

can you show me how to solve it once i subtract the 2x from both sides im stuck

OpenStudy (anonymous):

\[2x-5y=3\rightarrow-5y=3-2x \rightarrow5y=2x-3rightarrowy=(2/5)x-3/5\]

OpenStudy (anonymous):

finally y=(2/5)x-3/5

OpenStudy (anonymous):

from the general equation of the straight line y=mx+b m=2/5 b=-3/5

OpenStudy (anonymous):

thank you vwerymuch very much

OpenStudy (anonymous):

anytime :)